Six Measures of Solvency or Profitability

The following data were taken from the financial statements of Gates Inc. for the current fiscal year.

Property, plant, and equipment (net) $1,555,400
Liabilities:
Current liabilities $221,000
Note payable, 6%, due in 15 years 1,111,000
Total liabilities $1,332,000
Stockholders' equity:
Preferred $2 stock, $100 par (no change during year) $1,332,000
Common stock, $10 par (no change during year) 1,332,000
Retained earnings:
Balance, beginning of year $1,420,000
Net income 500,000 $1,920,000
Preferred dividends $26,640
Common dividends 117,360 144,000
Balance, end of year 1,776,000
Total stockholders' equity $4,440,000
Sales $30,388,500
Interest expense $66,660

Assuming that total assets were $5,483,000 at the beginning of the current fiscal year, determine the following. When required, round to one decimal place.

a. Ratio of fixed assets to long-term liabilities
b. Ratio of liabilities to stockholders' equity
c. Asset turnover
d. Return on total assets %
e. Return on stockholders’ equity %
f. Return on common stockholders' equity
Answer #1

Statement showing the computation of the ratios

Ratio of fixed assets to long-term liabilities 1.4:1
Ratio of liabilities to stockholders' equity 1:0.25
Asset turnover 8.63 Times
Return on total assets 14.21%
Return on stockholders’ equity 11.26%
Return on common stockholders' equity 37.54%

1) Ratio of fixed assets to long-term liabilities = Fixed assets/ Long-term Liabilties= $1555400/$1111000=1.4:1

2) Ratio of liabilities to stockholders' equity= liabilities/stockholders' equity = $1111000/$4440000= 1:0.25

3) Asset turnover= Net sales/ Average total assets = 30388500/3519200 = 8.63 Times

Average total assets= (5483000+1555400)/2= 3519200

4) Return on total assets = Net income/ Average total assets

= $500000/$3519200*100= 14.21%

5) Return on stockholders’ equity= Net income/stockholders' equity= $500000/$4440000*100= 11.26%

6)Return on common stockholders' equity= Net income/common stockholders' equity=$500000/$1332000= 37.54%